Q1) If the tax rate is 40 percent, compute the after-tax real interest rate in each of the following.

a) the nominal interest rate is 10 percent and the inflation rate is 5 percent.

b) the nominal interest rate is 6 percent and the inflation rate is 2 percent


Q2) Assume that the banking system has total initial checking deposits of GH¢ 100 billion. Assume also that required reserves are 10 percent of checking deposits, and that bank hold no excess reserves and household hold no currency.

a) (I) what is the money multiplier?

(ii) what is the money supply?


b) If the Central Bank now raises required reserves to 20 percent of deposits,

(I) what is the new multiplier?

(ii) what is the new money supply?

(III) what is the change in money supply from (a) to (b) in (GH¢)?


1
Expert's answer
2021-06-03T09:54:01-0400

Q1. After-tax real interest rate

a)     0.6(0.1)- 0.05= 1%

b)     0.6(0.06)- 0.02= 1.6%

Q2. Total reserve= GH¢ 100 billion

               Required Reserve (rr)= 10%

a)     

Money multiplier = 1/rr

= 1/0.1

= 10

Money supply= total reserve/ money multiplier

=100/10

= GH¢ 10 billion


b)     Rr = 20%

Money multiplier= 1/rr

=1/0.2

=5


Money Suplly= 100/5

= GH¢ 20 billion

 The change is 10 billion